A NEW VIRUS IS RELEASED ON THE INTERNET; THE ADMINISTRATOR OF A DEPARTMENTS LOCAL ARE NETWORK (LAN) IS GIVEN FIVE MINUTES BY A MANAGER TO ESTIMATE THE IMPACT. THE ADMINISTRATOR SAMPLES 12 OF THE PCs CONNECTED TO THE LAN, AND FINDS THAT 7 ARE INFECTED; USE PROPORTION TO ESTIMATE THE NUMBER OF INFECTED PCs IF THERE ARE A TOTAL OF 117 PCs CONNECTED TO THE LAN.
Answer by stanbon(75887)

A NEW VIRUS IS RELEASED ON THE INTERNET; THE ADMINISTRATOR OF A DEPARTMENTS LOCAL ARE NETWORK (LAN) IS GIVEN FIVE MINUTES BY A MANAGER TO ESTIMATE THE IMPACT. THE ADMINISTRATOR SAMPLES 12 OF THE PCs CONNECTED TO THE LAN, AND FINDS THAT 7 ARE INFECTED; USE PROPORTION TO ESTIMATE THE NUMBER OF INFECTED PCs IF THERE ARE A TOTAL OF 117 PCs CONNECTED TO THE LAN.
----------------
x/117=7/12
x = 117(7/12)
x is at least 69
===============
Cheers,
Stan H.